# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ IV.2 Quantum resource estimation studies for $p$-benzyne and FeMoco
In this section, we present QRE studies for two algorithmic approaches: the Trotter-based approach, and the qubitization-based double low-rank factorization algorithm originally proposed in Ref. von Burg et al. (2021a) and further analyzed in Ref. Lee et al. (2021). Moreover, for the Trotter-based algorithm, we report QRE results for two methods to ensure quantum computations within a target precision (for a discussion in greater detail, see Section A.2): the first approach is based on using rigorous analytic bounds on the errors resulting from the use of Trotter–Suzuki approximation and Trotterization, thus yielding a worst-case number of Trotter slices; the second approach relies on more-realistic Trotter numbers obtained through extrapolation from empirical studies of the Trotter–Suzuki errors for small circuits. In Sections A.1, A.2 and A.3, we elaborate on the workflow for generating the associated logical circuits and analyze the various bounds on the errors incurred in this process, as well as how we choose these bounds to ensure that quantum simulations achieve a given target accuracy.
@@
We report estimates for concrete physical resources required for a fault-tolerant implementation of the QPE algorithm for the $p$-benzyne and FeMoco molecules based on either the second-order Trotter–Suzuki formula used to approximate $\exp(-iHt)$ for the molecular Hamiltonian or the double-factorized (DF) qubitization algorithm of von Burg et al. von Burg et al. (2021a). In both cases, we assume access to a quantum state with significant overlap with the ground state as input to QPE, for example, a Hartree–Fock state. We do not include the cost of preparing this initial state in our resource estimations. It is worth emphasizing, however, that QPE is only provably fast for problems when the initial state is an eigenstate. When it is not an eigenstate of the Hamiltonian, there is a sampling overhead because the initial state is a superposition of eigenstates. This is a very problem-dependent challenge, but it can be ameliorated by classical preprocessing, that is, by running calculations on a classical computer to generate a better initial state which is then loaded into the quantum computer with a short quantum circuit. For example, one recent work Berry et al. (2024) presents an estimate that by using a matrix product state of bond dimension 4000, an overlap of 0.96 can be obtained for the ground state of the FeMoco molecule by implementing a circuit composed of nearly $10^{9}$ Toffoli gate.
@@
Here, $\epsilon_{\text{\tiny QPE}}$ is the desired error tolerance in phase estimation; $W(H)$ represents the unitary operator (as a function of the Hamiltonian $H$) used in the QPE algorithm (e.g., $W(H)=\exp(-iH\tau)$ in the case of Hamiltonian simulation for some time $\tau$, or $W(H)=e^{i\arccos(H/\lambda)}$ in the case of qubitization); $\Omega$ denotes the cost of a primitive circuit used to realize the implementation of $W(H)$ (such as a Trotter step in the Trotterization approach, or the LCU oracles associated with qubitization); and $g(\epsilon_{\text{\tiny QPE}})$ denotes the number of times that the primitive circuit must be repeated to ensure that the error in the spectrum of $H$ resulting from phase estimation of the eigenphases of $W(H)$ is at most $\epsilon_{\text{\tiny QPE}}$. Note that, in the qubitization approach, the operator $W(H)=e^{i\arccos(H/\lambda)}$ can typically be implemented as a quantum circuit exactly without approximations beyond those required for the synthesis of arbitrary-angle rotation gates; this implies $g(\epsilon_{\text{\tiny QPE}})=\mathcal{O}\left(1\right)$. Hence, due to $\|W^{\prime}(H)\|^{-1}\leq\lambda$, the overall cost of QPE in qubitization-based approaches becomes $\mathcal{O}\left(\Omega\lambda/\epsilon_{\text{\tiny QPE}}\right)$. Various versions of QPE have been analyzed in the literature, aiming to reduce its cost. For example, the standard QPE algorithm Nielsen and Chuang (2011) allows the estimation of eigenvalues within a target error $\epsilon_{\text{QPE}}$ with probability at least $1/2$ using $\lceil 16\pi/\epsilon_{\text{QPE}}\rceil$ applications of the unitary $\exp(-iHt)$. However, more optimized QPE strategies can achieve the multiplying factor (see Reiher et al. (2017); Babbush et al. (2018); Lee et al. (2021))
$M:=\lceil\pi/(2\epsilon_{\text{QPE}})\rceil$ (8)
@@
In Tables 3 and 4, we summarize our logical and physical QRE results for a number of circuits specified by the active space with sizes characterized by the number of orbitals $N_{\text{orb}}$, the number of logical qubits involved in the computation, and the overall allowable target error for QPE. To achieve chemical significance, the overall target error in ground-state energy estimation should be at least within “chemical accuracy”, that is, $\epsilon\leq 1.6\text{ mHa}$ (see, e.g., Ref. McArdle et al. (2020)). Energy estimations within chemical accuracy are often sufficient to predict important chemical properties such as chemical reaction rates, but even higher accuracies may be required for quantitatively precise predictions. Here we report resource estimates for two different precisions specified by either the allowable target error $\epsilon=1.0\text{ mHa}$ (for qualitative accuracy) or the much lower target error $\epsilon=0.1\text{ mHa}$ (for quantitative accuracy), using a circuit-level error budget of 0.01, respectively. The chemical basis set 6-31G is used to represent the spin orbitals in the case of $p$-benzyne, while for the FeMoco molecule we use the active-space model proposed by Li et al. Li et al. (2019).
In the case of the Trotter-based approach, physical runtimes for a complete implementation of QPE are obtained by multiplying the physical runtime for a single Trotter slice by the number of Trotter slices, and then by the number of controlled applications in QPE given by the value of $M$ in Equation 8. For the overall error budget of $\epsilon=0.1$ mHa, we obtain a value of $\epsilon_{\text{QPE}}=0.065$ mHa as an optimal choice (see section A.2), yielding $M=$ 24,167; for the overall target error $\epsilon=1.0$ mHa, we use $\epsilon_{\text{QPE}}=0.65$ mHa, yielding $M=$ 2417. The error budget allocation in the qubitization approach is discussed in Section A.3.
@@
As explained in Refs. Litinski (2019); Silva et al. (2024b, a), Clifford operations can be efficiently commuted to the end of the logical circuit by tracking a Clifford frame along the circuit. Some of the resulting non-Clifford gates can be merged into Clifford gates. Therefore, the process may be repeated until convergence. We call this procedure transpilation as explained in Section B.1. The outcome of transpilation is a sequence of non-Clifford gates in the form of multi-qubit $\pi/8$ Pauli rotations that must be executed using magic state injection. Therefore, the design of a fault-tolerant architecture that efficiently implements a given quantum algorithm reduces to constructing magic state factories (MSF) that can distill magic states of a target distance and fidelity at a rate on par with the fault-tolerant execution of non-Clifford gates. More details about the design of MSFs and the additional components of the layout are provided in Section B.2. In Table 4, we report the expected physical runtime and the number of physical qubits required for a fault-tolerant implementation of a quantum circuit when using hardware either with baseline or target parameter values, as well as for the $\Lambda_{18}$ noise model (representative of a desired hardware).
@@
Active space Number of orbitals, $N_{\text{orb}}$ $\epsilon=1.0$ mHa $\epsilon=0.1$ mHa
@@
$p$-benzyne, HL±$\pm 2$ 6 12 $9.5\times 10^{9}$ 12 $4.1\times 10^{11}$
$p$-benzyne, HL±$\pm 6$ 14 28 $9.5\times 10^{11}$ 28 $3.8\times 10^{13}$
$p$-benzyne, HL±$\pm 8$ 18 36 $4.1\times 10^{12}$ 36 $1.6\times 10^{14}$
$p$-benzyne, HL±$\pm 12$ 26 52 $3.6\times 10^{13}$ 52 $1.4\times 10^{15}$
@@
$p$-benzyne, HL±$\pm 2$ 6 12 $5.4\times 10^{8}$ 12 $2.5\times 10^{10}$
$p$-benzyne, HL±$\pm 6$ 14 28 $5.8\times 10^{9}$ 28 $4.0\times 10^{11}$
$p$-benzyne, HL±$\pm 8$ 18 36 $1.7\times 10^{10}$ 36 $8.0\times 10^{11}$
$p$-benzyne, HL±$\pm 12$ 26 52 $4.4\times 10^{10}$ 52 $2.3\times 10^{12}$
@@
$p$-benzyne, HL±$\pm 2$ 6 303 $7.8\times 10^{6}$ 341 $9.0\times 10^{8}$
$p$-benzyne, HL±$\pm 8$ 18 527 $3.9\times 10^{9}$ 568 $4.6\times 10^{10}$
$p$-benzyne, HL±$\pm 12$ 26 701 $1.5\times 10^{10}$ 748 $1.8\times 10^{11}$
FeMoco Li et al. (2019) 76 1792 $1.3\times 10^{12}$ 1972 $1.4\times 10^{13}$
@@
Baseline Parameter Set ($\bm{\Lambda\approx 2.34}$) Target Parameter Set ($\bm{\Lambda\approx 9.3}$) Desired Parameter Set ($\bm{\Lambda\approx 18}$)
@@
$1.0$ mHa 6 3.1$\times$107 5.3 days 15, 33, 79 | 87 1.6$\times$106 1.3 days 11, 29 | 33 6.6$\times$105 23.1 hours 23 | 25
14 3.6$\times$107 1.6 years 17, 37, 91 | 99 2.3$\times$106 142.2 days 13, 33 | 37 8.6$\times$105 111.5 days 25 | 29
18 3.8$\times$107 7.3 years 17, 39, 95 | 103 2.7$\times$106 1.8 years 13, 37 | 39 1.1$\times$106 1.3 years 27 | 29
26 5.0$\times$107 70.5 years 21, 39, 107 | 113 3.5$\times$106 15.5 years 15, 37 | 39 1.9$\times$106 12.3 years 11, 29 | 31
@@
$1.0$ mHa 6 1.1$\times$107 6.6 hours 29, 71 | 81 1.7$\times$106 1.5 hours 11, 27 | 29 5.5$\times$105 1.2 hours 21 | 23
14 3.5$\times$107 3.1 days 15, 33, 77 | 85 1.8$\times$106 17.4 hours 11, 29 | 31 7.1$\times$105 14.1 hours 23 | 25
18 3.4$\times$107 10.0 days 15, 35, 79 | 91 2.1$\times$106 2.5 days 11, 33 | 35 7.3$\times$105 1.8 days 23 | 25
26 3.3$\times$107 28.4 days 15, 35, 87 | 101 2.6$\times$106 5.9 days 13, 31 | 33 8.1$\times$105 4.8 days 23 | 27
@@
$1.0$ mHa 6 1.5$\times$107 5.4 minutes 25, 61 | 75 1.6$\times$106 1.2 minutes 23 | 27 8.9$\times$105 57.6 seconds 17 | 21
18 4.9$\times$107 2.2 days 15, 33, 75 | 91 4.0$\times$106 12.4 hours 11, 29 | 33 2.1$\times$106 10.2 hours 21 | 27
26 6.0$\times$107 10.0 days 15, 33, 87 | 103 5.5$\times$106 2.3 days 11, 31 | 37 2.8$\times$106 1.7 days 23 | 27
76 1.2$\times$108 2.6 years 17, 37, 93 | 109 1.5$\times$107 234.1 days 13, 33 | 43 8.0$\times$106 168.8 days 27 | 31
@@
We plot the results of our QRE studies, including both the runtime and the number of physical qubits, in Figure 29, alongside estimates for the runtime for two classical algorithms, the variational numerically exact full configuration interaction (FCI) computation and the heuristic density matrix renormalization group (DMRG) method White (1992), which were calculated by extrapolating the results of recent classical calculations Bellonzi et al. (2024); Gao et al. (2024). See Appendix D for details on this extrapolation.
(a) Qualitatively Accurate Simulation ($1.0$ mHa) (b) Quantitatively Accurate Simulation ($0.1$ mHa)
@@
This study demonstrates that ground-state energy estimation for molecules involving active spaces with orbital numbers in the range of 10 to 76 require a number of physical qubits ranging from approximately $10^{6}$ to $10^{8}$ and physical runtimes ranging from a few hours to several years. Both the quantum algorithm used and the hardware quality can have a significant impact on the resource requirements. On the algorithmic level, substantial space–time trade-offs can be observed. Implementations of the QPE algorithm based on Trotterization typically yield high $T$ counts resulting in long runtimes, while the required number of qubits to run the algorithm is low, whereas implementations based on qubitization result in much lower $T$ counts and higher qubit counts. For both algorithms, improving the hardware quality from baseline to target results in a reduced runtime and qubit count by approximately a factor of 5. Better algorithms run on better hardware can result in a reduction in runtime of up to two orders of magnitude. For example, an implementation of QPE with qubitization and target hardware results in runtime reduction by a factor of 50 compared to running QPE based on Trotterization (using empirical bounds) on baseline hardware. We also provide results using AzureQRE in Section B.3.
@@
Furthermore, we observe that, for $N_{\text{orb}}\gtrapprox 25$, quantum simulations begin to outperform classical FCI computations. Linear variational post-Hartree–Fock approaches based on the FCI method are designed to provide numerically exact solutions, but their practical use is known to be limited to molecular systems with few electrons and small basis sets. Although some molecular systems are classically tractable even at scales up to 100 orbitals, in general, exact classical computations become nearly impossible beyond 25 orbitals, especially for highly correlated systems. However, powerful classical heuristic algorithms can push the quantum advantage to much greater molecular sizes. For example, the DMRG method White (1992) run with parallel processing on HPC units can be significantly faster than quantum simulations for molecular active spaces involving up to $N_{\text{orb}}\approx 50$ spatial orbitals, as can also be observed in Figure 29(a). Nevertheless, DMRG methods eventually become increasingly unreliable for molecular systems involving a number of spatial orbitals far beyond 50, as such systems are typically too strongly correlated, requiring calculations with very large bond dimensions, and thus intractable runtimes McArdle et al. (2020); Reiher et al. (2017). While there is no such sharp transition line between what is classically tractable and intractable (which highly depends on the extent of quantum correlations in the studied systems), a quantum advantage gradually appears for orbital numbers beyond $N_{\text{orb}}\approx 50$. These insights also motivate future research, namely, developing quantum heuristic algorithms that could bring the transition to a quantum advantage down to smaller problem sizes. This naturally follows the development of classical algorithms, where the transition from the guarantees of FCI to the heuristics of DMRG greatly reduced the necessary resources.

### v2

@@ IV.2 Quantum resource estimation for $p$-benzyne and FeMoco
In this section, we present QRE studies for two algorithmic approaches: the Trotter-based approach, and the qubitization-based double low-rank factorization algorithm originally proposed in Ref. von Burg et al. (2021a) and further analyzed in Ref. Lee et al. (2021). Moreover, for the Trotter-based algorithm, we report QRE results for two methods to ensure quantum computations within a target precision (for a discussion in greater detail, see Section A.2): the first approach is based on using rigorous analytic bounds on the errors resulting from the use of Trotter–Suzuki approximation and Trotterization, thus yielding a worst-case number of Trotter slices; the second approach relies on more-realistic Trotter numbers obtained through extrapolation from empirical studies of the Trotter–Suzuki errors for small circuits. In Sections A.1, A.2 and A.3, we elaborate on the workflow for generating the associated logic circuits and analyze the various bounds on the errors incurred in this process, as well as how we choose these bounds to ensure that quantum simulations achieve a given target accuracy.
@@
We report estimates for concrete physical resources required for a fault-tolerant implementation of the QPE algorithm for the $p$-benzyne and FeMoco molecules based on either the second-order Trotter–Suzuki formula used to approximate $\exp(-iHt)$ for the molecular Hamiltonian or the double-factorized (DF) qubitization algorithm of von Burg et al. von Burg et al. (2021a). In both cases, we assume access to a quantum state with significant overlap with the ground state as input to QPE, for example, a Hartree–Fock state. We do not include the cost of preparing this initial state in our resource estimations. It is worth emphasizing, however, that QPE is only provably fast for problems when the initial state is an eigenstate. When it is not an eigenstate of the Hamiltonian, there is a sampling overhead because the initial state is a superposition of eigenstates. This challenge is problem-dependent but can be ameliorated by classical preprocessing, that is, by running calculations on a classical computer to generate a better initial state which is then loaded into the quantum computer with a short quantum circuit. For example, one recent work Berry et al. (2024) presents an estimate that by using a matrix product state of bond dimension 4000, an overlap of 0.96 can be obtained for the ground state of the FeMoco molecule by implementing a circuit composed of nearly $10^{9}$ Toffoli gate.
@@
Here, $\epsilon_{\text{\tiny QPE}}$ is the desired error tolerance in phase estimation; $W(H)$ represents the unitary operator (as a function of the Hamiltonian $H$) used in the QPE algorithm (e.g., $W(H)=\exp(-iH\tau)$ in the case of Hamiltonian simulation for some time $\tau$, or $W(H)=e^{i\arccos(H/\lambda)}$ in the case of qubitization); $\Omega$ denotes the cost of a primitive circuit used to realize the implementation of $W(H)$ (such as a Trotter step in the Trotterization approach, or the LCU oracles associated with qubitization); and $g(\epsilon_{\text{\tiny QPE}})$ denotes the number of times that the primitive circuit must be repeated to ensure that the error in the spectrum of $H$ resulting from phase estimation of the eigenphases of $W(H)$ is at most $\epsilon_{\text{\tiny QPE}}$. Note that, in the qubitization approach, the operator $W(H)=e^{i\arccos(H/\lambda)}$ can typically be implemented as a quantum circuit exactly without approximations beyond those required for the synthesis of arbitrary-angle rotation gates; this implies $g(\epsilon_{\text{\tiny QPE}})=\mathcal{O}\left(1\right)$. Hence, due to $\|W^{\prime}(H)\|^{-1}\leq\lambda$, the overall cost of QPE in qubitization-based approaches becomes $\mathcal{O}\left(\Omega\lambda/\epsilon_{\text{\tiny QPE}}\right)$. Various versions of QPE have been analyzed in the literature, aiming to reduce its cost. For example, the standard QPE algorithm Nielsen and Chuang (2011) allows the estimation of eigenvalues within a target error $\epsilon_{\text{QPE}}$ with probability at least $1/2$ using $\lceil 16\pi/\epsilon_{\text{QPE}}\rceil$ applications of the unitary $\exp(-iHt)$. However, more optimized QPE strategies can achieve the multiplying factor (see Refs. Reiher et al. (2017); Babbush et al. (2018); Lee et al. (2021))
$M:=\lceil\pi/(2\epsilon_{\text{QPE}})\rceil.$ (10)
@@
In Tables 4 and 5, we summarize our logical and physical QRE results for a number of circuits specified by the active space with sizes characterized by the number of orbitals $N_{\text{orb}}$, the number of logical qubits involved in the computation, and the overall allowable target error for QPE. To achieve chemical significance, the overall target error in ground-state energy estimation should be at least within “chemical accuracy”, that is, $\epsilon\leq 1.6\text{ mHa}$ (see, e.g., Ref. McArdle et al. (2020)). Energy estimations within chemical accuracy are often sufficient to predict important chemical properties such as chemical reaction rates, but even higher accuracies may be required for quantitatively precise predictions. Here we report resource estimates for two different precisions specified by either the allowable target error $\epsilon=1.6\text{ mHa}$ (for qualitative accuracy) or the much lower target error $\epsilon=0.1\text{ mHa}$ (for quantitative accuracy), using a circuit-level error budget of 0.01, respectively. The chemical basis set 6-31G is used to represent the spin orbitals in the case of $p$-benzyne, while for the FeMoco molecule we use the active-space model proposed by Li et al. Li et al. (2019).
In the case of the Trotter-based approach, physical runtimes for a complete implementation of QPE are obtained by multiplying the physical runtime for a single Trotter slice by the number of Trotter slices, and then by the number of controlled applications in QPE given by the value of $M$ in Equation 10. For the overall error budget of $\epsilon=0.1$ mHa, we obtain a value of $\epsilon_{\text{QPE}}=0.065$ mHa as an optimal choice (see Section A.2), yielding $M=$ 24,167; for the overall target error $\epsilon=1.6$ mHa, we use $\epsilon_{\text{QPE}}=1.04$ mHa, yielding $M=$ 1511. The error budget allocation in the qubitization approach is discussed in Section A.3.
@@
As explained in Refs. Litinski (2019); Silva et al. (2024b, a), Clifford operations can be efficiently commuted to the end of the logic circuit by tracking a Clifford frame along the circuit. Some of the resulting non-Clifford gates can be merged into Clifford gates. Therefore, the process may be repeated until convergence. We call this procedure transpilation as explained in Section B.1. The outcome of transpilation is a sequence of non-Clifford gates in the form of multi-qubit $\pi/8$ Pauli rotations that must be executed using magic state injection. Therefore, the design of a fault-tolerant architecture that efficiently implements a given quantum algorithm reduces to constructing magic state factories (MSF) that can distill magic states of a target distance and fidelity at a rate on par with the fault-tolerant execution of non-Clifford gates. More details about the design of MSFs and the additional components of the layout are provided in Section B.2. In Table 5, we report the expected physical runtime and the number of physical qubits required for a fault-tolerant implementation of a quantum circuit when using hardware with baseline, target, or desired parameter values.
@@
Active space Number of orbitals, $N_{\text{orb}}$ $\epsilon=1.6$ mHa $\epsilon=0.1$ mHa
@@
$p$-benzyne, HL±$\pm 2$ 6 12 $4.5\times 10^{9}$ 12 $4.1\times 10^{11}$
$p$-benzyne, HL±$\pm 6$ 14 28 $4.6\times 10^{11}$ 28 $3.8\times 10^{13}$
$p$-benzyne, HL±$\pm 8$ 18 36 $2.0\times 10^{12}$ 36 $1.6\times 10^{14}$
$p$-benzyne, HL±$\pm 12$ 26 52 $1.7\times 10^{13}$ 52 $1.4\times 10^{15}$
@@
$p$-benzyne, HL±$\pm 2$ 6 12 $2.7\times 10^{8}$ 12 $2.5\times 10^{10}$
$p$-benzyne, HL±$\pm 6$ 14 28 $3.5\times 10^{9}$ 28 $4.0\times 10^{11}$
$p$-benzyne, HL±$\pm 8$ 18 36 $1.1\times 10^{10}$ 36 $8.0\times 10^{11}$
$p$-benzyne, HL±$\pm 12$ 26 52 $2.7\times 10^{10}$ 52 $2.3\times 10^{12}$
@@
$p$-benzyne, HL±$\pm 2$ 6 291 $4.9\times 10^{6}$ 341 $9.0\times 10^{8}$
$p$-benzyne, HL±$\pm 8$ 18 502 $2.3\times 10^{9}$ 568 $4.6\times 10^{10}$
$p$-benzyne, HL±$\pm 12$ 26 668 $9.0\times 10^{9}$ 748 $1.8\times 10^{11}$
FeMoco Li et al. (2019) 76 1789 $7.7\times 10^{11}$ 1972 $1.4\times 10^{13}$
@@
Baseline Parameter Set Target Parameter Set Desired Parameter Set
@@
$1.6$ mHa 6 3.5$\times$107 2.5 days 15, 33, 75 | 87 1.6$\times$106 13.7 hours 11, 29 | 31 4.4$\times$105 11.1 hours 21 | 25
14 3.7$\times$107 282.8 days 17, 37, 87 | 97 2.2$\times$106 64.9 days 13, 33 | 35 7.6$\times$105 50.1 days 25 | 27
18 3.9$\times$107 3.5 years 17, 39, 91 | 101 2.4$\times$106 293.7 days 13, 35 | 37 9.2$\times$105 230.2 days 27 | 29
26 4.5$\times$107 33.5 years 17, 43, 97 | 111 3.2$\times$106 7.5 years 15, 37 | 39 1.1$\times$106 6.0 years 29 | 31
@@
$1.6$ mHa 6 1.1$\times$107 3.2 hours 29, 69 | 77 6.7$\times$105 46.5 minutes 25 | 29 3.7$\times$105 36.9 minutes 19 | 23
14 4.1$\times$107 2.1 days 19, 31, 75 | 95 1.6$\times$106 10.7 hours 11, 29 | 31 4.9$\times$105 8.6 hours 21 | 25
18 3.1$\times$107 6.4 days 15, 33, 79 | 95 1.5$\times$106 1.5 days 11, 29 | 35 6.3$\times$105 1.1 days 23 | 25
26 3.7$\times$107 16.0 days 15, 35, 81 | 93 2.0$\times$106 3.8 days 13, 29 | 33 6.7$\times$105 2.7 days 23 | 25
@@
$1.6$ mHa 6 1.4$\times$107 3.3 minutes 25, 59 | 73 1.3$\times$106 1.2 minutes 23 | 27 8.2$\times$105 35.8 seconds 17 | 21
18 4.8$\times$107 1.4 days 15, 31, 75 | 93 3.6$\times$106 7.3 hours 11, 27 | 33 1.7$\times$106 5.6 hours 21 | 25
26 5.5$\times$107 5.4 days 15, 33, 79 | 95 4.7$\times$106 1.3 days 11, 29 | 35 2.6$\times$106 23.6 hours 23 | 27
76 1.2$\times$108 1.4 years 17, 37, 89 | 107 1.3$\times$107 121.8 days 13, 33 | 39 7.7$\times$106 96.8 days 25 | 31
@@
We plot the results of our QRE studies, including both the runtime and the number of physical qubits, in Figure 30, alongside estimates for the runtime for two classical algorithms, the variational numerically exact full configuration interaction (FCI) computation and the heuristic density matrix renormalization group (DMRG) method White (1992), which were calculated by extrapolating the results of recent classical calculations Bellonzi et al. (2024a); Gao et al. (2024). See Appendix D for details on this extrapolation.
(a) Qualitatively Accurate Simulation ($1.6$ mHa) (b) Quantitatively Accurate Simulation ($0.1$ mHa)
@@
This study demonstrates that ground-state energy estimation for molecules involving active spaces with orbital numbers in the range of 10 to 76 require a number of physical qubits ranging from approximately $10^{6}$ to $10^{8}$ and physical runtimes ranging from a few hours to several years. Both the quantum algorithm used and the hardware quality can have a significant impact on the resource requirements. On the algorithmic level, substantial space–time trade-offs can be observed. Implementations of the QPE algorithm based on Trotterization typically yield high $T$ counts resulting in long runtimes, while the required number of qubits to run the algorithm is low, whereas implementations based on qubitization result in much lower $T$ counts and higher qubit counts. For both algorithms, improving the hardware quality from baseline to target results in a reduced runtime and qubit count by up to a factor of 5. Better algorithms run on better hardware can result in a reduction in runtime of up to two orders of magnitude. For example, an implementation of QPE with qubitization and target hardware results in runtime reduction by a factor of 50 compared to running QPE based on Trotterization (using empirical bounds) on baseline hardware. We also provide results using AzureQRE in Section B.3.
@@
Furthermore, we observe that, for $N_{\text{orb}}\gtrapprox 25$, quantum simulations begin to outperform classical FCI computations. Linear variational post-Hartree–Fock approaches based on the FCI method are designed to provide numerically exact solutions, but their practical use is known to be limited to molecular systems with few electrons and small basis sets. Although some molecular systems are classically tractable even at scales up to 100 orbitals, in general, exact classical computations become nearly impossible beyond 25 orbitals, especially for highly correlated systems. However, powerful classical heuristic algorithms can push the quantum advantage to much greater molecular sizes. For example, the DMRG method White (1992) run with parallel processing on HPC units can be significantly faster than quantum simulations for molecular active spaces involving up to $N_{\text{orb}}\approx 50$ spatial orbitals, as can also be observed in Figure 30(a). The most recent cutting-edge petaFLOPS performance results for a hybrid CPU–multi-GPU parallel implementation of the spin-adapted ab initio DMRG method on the state-of-the-art Nvidia DGX-H100 architecture suggest that heuristic classical simulations of molecular systems are tractable even for sizes well above 75 spatial orbitals Menczer et al. (2024). Nevertheless, DMRG methods eventually become increasingly unreliable for molecular systems involving a number of spatial orbitals far beyond 50, as such systems are typically too strongly correlated, requiring calculations with large bond dimensions, and thus intractable runtimes McArdle et al. (2020); Reiher et al. (2017). While there is no such sharp transition line between what is classically tractable and classically intractable (which is highly dependent on the extent of quantum correlations in the studied systems), a quantum advantage gradually appears for spatial orbital numbers beyond $N_{\text{orb}}\approx 50$. These insights also motivate future research, namely, developing quantum heuristic algorithms that could bring the transition to a quantum advantage down to smaller problem sizes. This naturally follows the development of classical algorithms, where the transition from the guarantees of FCI to the heuristics of DMRG greatly reduced the necessary resources.
[…]
It is insightful to understand the discrepancy between the runtimes reported here for the DF qubitization algorithm and the four days reported for THC qubitization in Ref. Lee et al. (2021), for quantum simulations of the Li et al. FeMoco within the chemical accuracy 1.6 mHa. For instance, the 122 days and 97 days that we report as runtimes for the target and desired hardware specifications, respectively (see Table 5), are longer by the factors of approximately 30 and 24 than the four days reported in Ref. Lee et al. (2021). We note that, on the logical level, the $T$-gate count reported here for DF qubitization (see Table 4) is higher by only a factor of 6 than the $T$ gate count resulting from converting the Toffoli gate count reported for THC qubitization in Table III of Ref. Lee et al. (2021), assuming that at least four $T$ gates are required for decomposing each Toffoli gate into gates from the Clifford+$T$ gate set; importantly, apart from using different representations of the Hamiltonian in the two algorithms (both based on qubitization), the use of different criteria for determining the truncation thresholds in the double factorization and tensor hypercontraction approaches can significantly contribute to the difference in the reported $T$-gate counts. Thus, the discrepancy between the reported runtimes is not entirely due to running a different quantum simulation algorithm (DF qubitization vs. THC qubitization), but also to differing assumptions on the physical and architectural levels. The resource requirements reported in Ref. Lee et al. (2021) are based on using AutoCCZ states (and the associated CCZ factories) to implement the Toffoli gates, while the resource estimates reported here are based on $T$ magic states (and the associated factories) to implement the $\pi/8$ rotations. We expect that supporting CCZ factories in our architecture will allow for up to a four times faster consumption of magic states depending on the input logic circuit. Together, the factor of 4 (on the architectural level) and factor of 6 (on the logical level) result in a factor of 24 for the expected difference between the runtime for DF qubitization reported in this work and the runtime for THC qubitization in Lee et al. (2021), which explains the difference in 122 days (or 97 days) versus four days. Further differences arise due to using different hardware specifications. The physical noise model assumed in Ref. Lee et al. (2021) results in the values $\Lambda\approx 6.1$ and $\bar{\Lambda}\approx 5.0$ for the error suppression rates associated with the error suppression models discussed in Section III.2. In addition, different physical space–time trade-offs can even be obtained for the same logic circuit and the same noise model, as is illustrated in Figure 31 for the circuit associated with quantum simulation of the Li et al. FeMoco based on THC qubitization, given the architecture discussed in this paper. Such space–time trade-off considerations are discussed in more detail in Ref. Silva et al. (2024a).
[…]